### v2

# WildLive: Near Real-time Visual Wildlife Tracking onboard UAVs

Abstract
Live tracking of wildlife via high-resolution video processing directly onboard drones is widely unexplored and most existing solutions rely on streaming video to ground stations to support navigation. Yet, both autonomous animal-reactive flight control beyond visual line of sight and/or mission-specific individual and behaviour recognition tasks rely to some degree on this capability. In response, we introduce WildLive – a near real-time animal detection and tracking framework for high-resolution imagery running directly onboard uncrewed aerial vehicles (UAVs). The system performs multi-animal detection and tracking at 17fps+ for HD and 7fps+ on 4K video streams suitable for operation during higher altitude flights to minimise animal disturbance. Our system is optimised for Jetson Orin AGX onboard hardware. It integrates the efficiency of sparse optical flow tracking and mission-specific sampling with device-optimised and proven YOLO-driven object detection and segmentation techniques. Essentially, computational resource is focused onto spatio-temporal regions of high uncertainty to significantly improve UAV processing speeds without domain-specific loss of accuracy. Alongside, we introduce our WildLive dataset, which comprises 200k+ annotated animal instances across 19k+ frames from 4K UAV videos collected at the Ol Pejeta Conservancy in Kenya. All frames contain ground truth bounding boxes, segmentation masks, as well as individual tracklets and tracking point trajectories. We compare our system against current object tracking approaches including OC-SORT, ByteTrack, and SORT. Our multi-animal tracking experiments with onboard hardware confirm that near real-time high-resolution wildlife tracking is possible on UAVs whilst maintaining high accuracy levels as needed for future navigational and mission-specific animal-centric operational autonomy. We publish all source code, weights, dataset, and labels for easy utilisation by the community.
1 Introduction and Motivation
Live Tracking of Animals via Drones. Multi-object tracking (MOT) [9, 35] in high-resolution video streams processed live onboard drones [28] poses significant challenges when applied to wildlife monitoring [29, 31, 10, 22] due to environmental demands, small animal resolution [33], platform motion, computational constraints, energy limitations and more. Yet, this capability plays a key role in enabling future animal-reactive and/or Beyond Visual Line of Sight (BVLOS) flight control for mission-specific individual [2, 25] and behaviour recognition [6, 5, 21, 8] – without options to involve ground control due to latency or connectivity constraints. Although first attempts [24] to build such systems exist, operation is so far limited to low resolutions without full benchmarking where datasets/code are so far not fully public111Main parts of the source code for [24] is available at https://github.com/hardboy12/YOLOv7-DeepSORT, yet the repository is incomplete. Contacting the author re sharing full source materials and their dataset for reproducibility/benchmarking purposes did not result in a response before the date of this paper’s submission.. In response, this paper proposes, benchmarks and shares with the community the near real-time MOT WildLive system (see Fig. 1) suitable for advanced tracking of animals in high-resolution video streams during flight directly onboard UAVs with a Jetson Orin AGX.
Our main contributions in this work are:
1.
We introduce and make publicly available222Our materials can be found at https://dat-nguyenvn.github.io/WildLive/ our MOT WildLive system (see Sec. 4) for near real-time wildlife tracking optimised for and deployable directly on drones that carry an embedded Jetson Orin AGX computer.
2.
We benchmark WildLive against suitable SOTA systems on a domain-specific tracking dataset (see Sec. 3) which we introduce alongside our system. Amongst other rich annotations, it contains 215K+ tracked bounding boxes from representative, UAV-acquired video sequences recorded on site in Kenya under strict ethical oversight (see Ethics Statement).
3.
We publish2 our WildLive Benchmark Dataset and its ground truth information in full for reproducibility and domain-relevant comparability in this evolving field.
2 Paper Concept and Related Work
Detection vs. Tracking. Whilst localising the presence of objects in video, i.e. detection, requires matching complex pixel patterns over potentially large object regions, following content across frames, i.e. tracking, may either utilise these entire objects [34, 11, 3, 9] or alternatively follow discrete, potentially sparse locations on the objects [12, 13, 19] only. ByteTrack [34], OCSORT [7], and SORT [3] are examples of recent full object trackers computationally suitable for edge device deployment. On the other hand, deep trackers such as CoTracker [19] and traditional sparse optical flow trackers such as Lucas-Kanade (LK) [4] implement point tracking capabilities. The latter are vastly cheaper computationally, but have shortcomings regarding occlusions, aperture limitations and viewpoint changes. Given that object (re)detection may not be required every frame, this offers performance headroom to combine and balance fast LK tracking with intermittent deep detection to perform light-weight, UAV-suitable wildlife tracking in high-resolution video near real-time.
Slicing Aided Hyper Inference (SAHI) and YOLO. To date, the YOLO [30, 23, 32, 17, 20] detector series have remained amongst the fastest deep object detectors of their time throughout their version history. To process high-resolution 4K images with YOLO without performance-crushing information loss due to downsampling, we apply the SAHI technique [1] in conjunction with YOLOv8 [17] or YOLO11 [20] for optimised processing speeds beyond simple window processing without loss of accuracy. Note that all performance metrics are evaluated under this same regime including tracker benchmarking for a fair comparison and maximal utilisation of video content – particularly given small animal sizes in most UAV-acquired footage.
3 Dataset
The WildLive Benchmark Dataset. Our dataset contains 215,800 bounding boxes and animal segmentation masks along its 291 zebra, giraffe, and elephant tracklets, plus 84 point tracks across 22 UAV-acquired 4K video sequences, totaling 19,139 frames recorded on site at the Ol Pejeta Conservancy in Kenya. Acquisition was conducted via DJI Mavic 3 Enterprise and Pro drones plus a custom-built quadcopter for wildlife missions. Figs. 2 and 3 exemplify frames and key metadata. Overall, the dataset provides verified animal bounding boxes and SAM2 [26] segmentations, tracklet IDs, as well as manually corrected, sparse LK pixel trajectories.
4 Method
Framework Overview. As summarised in Fig. 1, our WildLive framework is optimised for near real-time high-resolution video stream tracking onboard drones. It integrates SAHI sampling and YOLO object detection with light-weight sparse optical flow LK tracking and YOLO instance segmentation to localise and follow animals live on UAV platforms. The following sections describe the framework and provide technical details on our system.
4.1 Detection, Segmentation and Point Selection
Initialisation and Strategic Sampling. High-resolution frames are first processed via SAHI [1] using YOLO11 [20] as detection (and during tracking intialisation also segmentation) model to provide fast localisation of low-resolved animals. This full frame scan at is computationally expensive and cannot operate near real-time at high resolutions given today’s UAV hardware. Thus, it is used to initiate tracking, whilst later detection and corroboration only focuses on localized 640640 windows with highest strategic update need. Biological systems focus computational resource in similar ways via foveal vision [27, 14]. After initiation, two local region categories are prioritised above others for more frequent detection window probing:
Frame Edges: to detect new animals entering the frame.
Tracked Instances: to validate current tracks and to detect disappearances or occlusions.
Thus, re-detections are applied to all image windows cyclically, however, at higher temporal sampling rate for above frame regions to focus computational resource.
Sparse Location Selection. Selectively processing only point locations on animals rather than performing dense, object-wide, or even frame-wide dense tracking provides order-of-magnitude faster performance to track and thereby interpolate between re-detections. points at locations per animal ID are derived as classic Harris corners [16] within the animal segment to be tracked where provided reliable performance. We note that with the system is not sensitive to small changes in . During re-detection validation of tracked animals, these tracked points are resampled to avoid drift, but are required to widely remain within the predicted YOLO segmentation mask to confirm object persistence (see source code for full details).
4.2 Integration with LK Tracking
Optimised Pyramidal LK Tracker. Expanding on early ideas for detector and LK tracker integration [6] and avoiding frame-by-frame filtered integration for instance via a Kalman Filter [15, 18], we use the pyramidal Lucas-Kanade (LKP) [4] only as a sparse interpolator between intermittent re-detections to boost speed under a high-resolution regime. Pyramidal processing of the vector of all tracked points from the coarsest level to the finest resolution level at logarithmic resolution scaling according to
can effectively support rapid high-resolution LKP processing of 4K+ imagery. Empirically, a value of yielded fastest overall results without loss of accuracy for our dataset (see Tab. 2).
Linking Point Tracking to Animal Tracklets. After determining LK displacement vectors for tracked points, full animal bounding box shifts are predicted as average displacement of the points associated with an animal instance. Noting that individual LKP tracks may fail due to occlusions, aperture limitations, viewpoint changes and more, model drift is possible. To address this, regular re-detections and point re-initialisations on the animal segments are performed, effectively using LKP tracking as a data-driven interpolator with minute computational footprint. Performing re-detection updates requires rapid matching of track IDs to re-detected animal masks.
Tracklet Lifecycle and Confidence. Let a Point-over-Area (PoA) index be defined as the proportion of tracked points with whose locations fall within a segmentation mask. Matching re-detections and point clouds associated to animal tracklets based on this index together with Intersection over Union (IoU) considerations to address overlap scenarios provides rapid re-association capability between LKP point propagations and YOLO re-detections (see source code for implementation detail). As a result, each YOLO re-detected object is either assigned to an existing animal tracklet or tracked forwards with a new ID if it does not sufficiently match any existing ID. Following [6], every tracklet also carries a confidence measure, where in our WildLive system, the measure itself accumulates YOLO re-detection confidence values minus a minimal required confidence per detection over time (see [6] for full details). This implements a simple and effective temporal evidence accumulator which, via basic thresholding, controls tracklet termination as well as tracklet validation, that is accepting a tracklet as ’1-confident’ (otherwise ’0-spurious’) and labelling it as such to the user. The latter allows the system to track even ’spurious’ YOLO detection candidates with low confidence in order to focus computational resources to validate or dismiss those, whilst only ’confident’ tracklets are considered for experimentation.
fps | MOTA | IDF1 | |||
| ByteTrack[34] (2022) + SAHI[1] | 8x | 0.33 | 65.34 | 60.59 | |||
| 8l | 0.41 | 65.67 | 63.05 | ||||
| 8m | 0.53 | 63.92 | 58.24 | ||||
| 8s | 0.69 | 62.55 | 57.28 | ||||
| 8n | 0.74 | 62.92 | 57.02 | ||||
| 11x | 0.31 | 72.50 | 68.44 | ||||
| 11l | 0.42 | 67.36 | 61.91 | ||||
| 11m | 0.50 | 66.62 | 60.64 | ||||
| 11s | 0.62 | 66.52 | 60.70 | ||||
| 11n | 0.67 | 62.43 | 55.45 | ||||
| OC-SORT[7] (2023) + SAHI[1] | 8x | 0.33 | 67.98 | 65.96 | |||
| 8l | 0.43 | 67.69 | 66.35 | ||||
| 8m | 0.55 | 62.11 | 59.35 | ||||
| 8s | 0.74 | 53.38 | 54.35 | ||||
| 8n | 0.79 | 47.62 | 48.15 | ||||
| 11x | 0.31 | 70.82 | 67.90 | ||||
| 11l | 0.44 | 64.19 | 63.63 | ||||
| 11m | 0.52 | 62.94 | 61.00 | ||||
| 11s | 0.66 | 56.90 | 55.75 | ||||
| 11n | 0.71 | 50.29 | 49.90 | ||||
| SORT[3] (2016) + SAHI[1] | 8x | 0.33 | 74.41 | 55.69 | |||
| 8l | 0.43 | 74.77 | 57.55 | ||||
| 8m | 0.56 | 68.28 | 51.23 | ||||
| 8s | 0.74 | 61.12 | 48.89 | ||||
| 8n | 0.79 | 57.93 | 44.14 | ||||
| 11x | 0.31 | 75.83 | 60.12 | ||||
| 11l | 0.46 | 71.08 | 55.60 | ||||
| 11m | 0.51 | 68.47 | 56.10 | ||||
| 11s | 0.66 | 66.31 | 50.31 | ||||
| 11n | 0.71 | 59.60 | 45.91 | ||||
| WildLive (Ours) | 8x | 4.79 | 76.65 | 75.86 | |||
| 8l | 5.29 | 78.70 | 79.03 | ||||
| 8m | 5.60 | 75.51 | 74.18 | ||||
[…]
| 8n | 5.68 | 70.29 | 69.31 | ||||
[…]
| 11m | 5.30 | 77.74 | 78.71 | ||||
| 11s | 5.76 | 75.52 | 74.76 | ||||
| 11n | 6.32 | 75.07 | 73.45 |
5 Experiments
Experimental Setup. We evaluate WildLive on offline Tesla P100-PCIE-16GB GPU hardware (see Fig. 1) and pinpoint its speed on the Jetson Orin AGX onboard GPU environment (see Fig. 2). We utilise standard networks (e.g. YOLO, SAHI as cited) for Tesla experiments, but based on these perform TensorRT optimisation of networks for maximum efficiency during Jetson experiments sharing2 all optimised network weights. Jetson deployment is further facilitated via Docker containerisation of the system offering deployment flexibility and scalability for the community. Successful test flights in Kenya (shown in Fig. 1) constitute a physical proof-of-concept that the full WildLive system can indeed operate on a custom built, flight-tested UAV2.
6 Results
Comparative System Benchmarks. We compare WildLive against recent SOTA full object trackers computationally suitable for edge device deployment, in particular ByteTrack [34], OCSORT [7], and SORT [3]. We benchmark processing speeds together with both Multi Object Tracking Accuracy (MOTA) and Identification F1 (IDF1) scores. The latter complements MOTA regarding limitations on how long trackers correctly identify objects. Full results are shown in Tab. 1 and confirm order-of-magnitude gains in processing speed (up to 6.32fps) compared to other tested techniques resting on only intermittent re-detection bridged by computationally negligible LKP point tracking. Accuracy, maybe surprisingly, slightly improves too, leading to best MOTA at and IDF1 at : utilising PoA and IoU together for ambiguity resolution proves superior compared to IoU-centred full object tracking methods tested which have no direct access to segmentation masks.
Runtine Speed Estimation. In order to estimate performance on drone hardware specifically, we also benchmark WildLive’s TensorRT-optimised processing speed on Jetson hardware as shown in Tab. 2 for a range of re-detection regimes scaling from our standard system (single window re-detection per frame named Ours) at 7.53fps down to the base case of permanent full frame re-detection at 2.45fps. We finally note that Ours processes HD streams at 17fps+.
7 Conclusion and Future Work
We conclude that WildLive’s high-resolution, high-speed tracking capabilities are close to those needed to fuel seamless, animal-reactive drone navigation based on video processing directly onboard UAVs, albeit accuracies under a small object regime still leave significant room for detection and tracking improvements. Such availability could revolutionise the way UAV conservation missions are conducted – in terms of range, cost, ease of use as well as mission type. We are in the process of producing full speed-accuracy benchmarks for multiple runs across the parameter space of WildLive and envisage rigorous field testing of the prototype next. Eventually, we will link our computer vision methods directly to UAV navigation for vision-driven autonomous BVLOS wildlife monitoring missions.
| 01 (Standard System (Ours)) | 7.53 | ||||
[…]
| 04 | 5.87 | ||||
| 08 | 4.59 | ||||
| 16 | 3.28 | ||||
| 24 (Full Frame Re-detection) | 2.45 |
Ethics Statement
When collecting our WildLive Benchmark Dataset and during test flights, we adhered to strict ethical standards to ensure the wellbeing of the animals and the appropriateness and integrity of the research. Our data was collected using drones in the Ol Pejeta Conservancy, Kenya. Drones were operated at safe distances to avoid causing stress or disturbance to animals. Additionally, all recordings were made in a non-invasive manner, with no direct interaction with wildlife. To further protect privacy and adhere to ethical guidelines, we ensured that human faces are not contained in recordings, focusing solely on animals and their behavior. All data collection procedures followed relevant local regulations, such as those provided by the Kenya Civil Aviation Authority (KCAA), the Kenya Wildlife Research and Training Institute (KWRTI), and other ethical guidelines. Efforts were made to respect the natural behaviours of the animals maximally. By prioritising animal welfare, privacy, and ethical research practices, the dataset can contribute to scientific advancements while minimising any harm to the environment and its inhabitants.